# Image Generation Models: A Technical History

### 08 Mar 2026 v1

@@ 7 Recent Developments in Image Generation
Many top image generation systems have shifted from classic diffusion based formulations and moved toward Rectified Flow and Flow Matching, motivated by cleaner continuous time dynamics, improved training stability at scale, image generation in fewer sampling steps. Moreover, the overall system has become more complicated and now consists of many pieces, including a large language model providing strong text understanding and better instruction following, more capable text encoders for precise conditioning, and a strong image backbone (often a Transformer) that converts those rich semantics into visual structure and detail. In the remainder of this section, we focus on the modeling shift itself: Rectified Flow and Flow Matching. Recently, they have emerged as compelling alternatives to diffusion objectives. We review the core formulation, and the key training and sampling choices that make these methods work well in practice. We also highlight the most impactful recent results that demonstrate their advantages for high quality and efficient image generation.
@@
$x_{t+1}=x_{t}+f\!\left(x_{t},\theta_{t}\right),$ (101)
@@ 7.1 Rectified Flow
$\frac{dZ_{t}}{dt}=v_{\theta}\!\bigl(Z_{t},t\bigr),$ (105)
@@
Later, Esser et al. [113] scaled Rectified Flow for high-resolution image generation and were able to achieve superior performance compared to established diffusion models. While their training objective follows Rectified Flow, for the text to image architecture they use a multimodal diffusion backbone, DiT [102]. Fig. 70 shows the overall architecture of their model.
@@
With the Rectified Flow training setup, they are able to achieve competitive results and outperform the best open-source models and DALLĒ 3 on GenEval [114]. Fig. 71 shows some of the sample images generated by their model with the corresponding prompts.
@@ 7.2 Flow Matching
with boundary conditions $\mu_{0}=0$, $\sigma_{0}=1$ at $t=0$ and $\mu_{1}=x_{1}$, $\sigma_{1}=\sigma_{\min}$ at $t=1$. They define an affine conditional map $\psi_{t}(\cdot)$ so that it pushes a standard Gaussian to the Gaussian conditional $p_{t}(x\mid x_{1})$
@@
where $\frac{d}{dt}\psi_{t}(z_{0};x_{1})$ is computed analytically from the definition of $\psi_{t}$. The authors show that, depending on the choice of the Gaussian path (i.e., $\mu_{t},\sigma_{t}$), the resulting probability path can correspond to either a diffusion path or an OT path. The simplified training and inference algorithm for FM is shown in Box 11
They compare results using the same model architecture but trained with different objectives: one trained with OT via Flow Matching, and another trained with a diffusion objective (DDPM [83]). As shown in Table XI, Flow Matching beats the diffusion based model in likelihood (BPD), quality of generated samples (FID), and evaluation time (NFE).
@@ 7.3 Conclusion
Box 11 A simplified view of Flow Matching training (via regression to a target velocity along a chosen conditional path) and inference (via an Euler sampler).
@@ 8.1 Video Generation with GANs
Figure 73: Sample generated video frames showing people walking on a golf course. Image source: [119]
@@
$\hat{x}^{(k)}_{I}=G_{I}\left(z_{I}^{(k)}\right),\qquad\hat{x}^{(k)}_{I}\in\mathbb{R}^{64\times 64\times 3}.$ (116)
@@
The $D_{I}$ terms are the losses for the image discriminator, and the $D_{V}$ terms are the losses for the video discriminator. Using this approach, they produce video outputs that score a much higher preference by the user, as shown in Table XII.
@@ 8.3 Diffusion Based Video Generation
Using this approach they are able to dramatically increase the inception score for unconditional video generation on the UCF-101 dataset, as shown in Table XIII. Fig. 77 shows text-conditioned video samples by cascading two models. First, samples are generated using a lower-resolution $16\times 64\times 64$ model. Then the samples are passed to a second model for simultaneous super-resolution and autoregressive extension to $64\times 128\times 128$.
@@ 8.4 Scaling up Video Diffusion Models
Blattmann et al. [131] introduced one of the first widely used open-source latent video diffusion models and called it Stable Video Diffusion (SVD). Most works prior to SVD only trained the video generation models on a small video dataset. Moreover, prior works mostly focused on temporal architectural changes rather than data and training strategy. The training pipeline consists of three stages. In stage I, they perform image pretraining with text–image pairs. Their experiments show that humans strongly prefer image pretrained models. In stage II, they pretrain the base video model. They preprocess a video clip and select $T$ frames and resize each frame to $H\times W=320\times 576$. They build conditioning by passing the text caption of the input video to a text encoder. The frame rate information is also embedded into a vector. The conditioning embeddings are then injected into the UNet block. The UNet block consists of both spatial (height, width) and temporal (frame) attention layers. They utilize the EDM [100] loss for training the UNet backbone.
Stage III has the same pipeline as stage II but they use 1 million high-quality text-to-video pairs that consist of lots of object motion, clean steady camera paths, and good captions. The resolution for this stage is $576\times 1024$ compared to $320\times 576$ at stage II. Box 12 shows their overall training pipeline. The authors also train an image-to-video variant of the model. For those, they start from the same base UNet and replace the text embedding with CLIP image embeddings and a few more tweaks.
@@
Fig. 78 shows some of the sample video frames generated by Stable Video Diffusion for different input text prompts.
@@ 9.4 DeepFake Detection and Benchmarks
Nataraj et al. [155] shifted towards an end-to-end framework to detect GAN-generated images, as shown in Fig. 90. They no longer rely on handcrafted features and instead let the CNN learn discriminative patterns from co-occurrence matrices computed separately on the R, G, and B channels.
They evaluated the performance of the model on two different datasets: CycleGAN [162] and StarGAN [163]. CycleGAN contains image-to-image style transfer samples, and StarGAN consists of images of real celebrities, as well as GAN-generated celebrity images. In order to evaluate how well the model generalizes, they train on one dataset and use the other dataset for testing. Table XV shows the performance of the model on the CycleGAN and StarGAN datasets.
@@ 9.6 Watermarking Generated Images
As image generation models improved with the introduction of diffusion-based models, it became increasinlgy difficult to detect images generated by diffusion models. Wang et al. [166] ask what happens if they take the image through the inversion and reconstruction path. For diffusion-generated images, there should be a noise path that leads to a better reconstruction. In order to showcase this, they first take the image through the DDIM inversion process:
@@
$\mathrm{DIRE}(x_{0})=\lVert x_{0}-x_{0}^{\prime}\rVert,$ (132)
where $x_{0}$ is the original image and $x_{0}^{\prime}$ is the reconstructed image.
Fig. 94 shows the DIRE representation of real and diffusion-generated images. In experiments, real images show larger DIRE values compared to diffusion-generated ones. Lastly, they train a binary (ResNet-based) classifier on DIRE to detect diffusion-generated images. Using this approach, they are able to achieve better accuracy and average precision compared to other state-of-the-art detector models.
@@
where $\lambda_{i}$ is a weight on the image similarity loss.
Note that only the signed decoder parameters of $D_{m}$ are updated but the encoder $E$ and extractor $W$ are kept frozen. Using a 48-bit signature, they are able to detect $99\%$ of newly generated images on the MS-COCO dataset while falsely flagging real images as fake only about one in $10^{9}$. Compared to Corvi et al. [168], they achieve an order of magnitude lower false positive rate for a similar true positive rate.
@@ 9.7 Conclusion
Image generation models have made significant leaps over the past decade. We have moved from models that produced low-quality, smudged outputs to models that generate images and coherent videos that are nearly indistinguishable from reality. Recent empirical scaling trends suggest continued improvements the capabilities of these models. A central challenge is establishing effective safeguards against potential misuses. Do we have adequate safety measures in place? Are we prepared for a future in which the harmful uses of image generation systems, described above, are realized at scale? Given the risks associated with these powerful image and video generation models, we need both technical and societal solutions to mitigate these risks and to ensure preparedness for worst case scenarios.

### 29 Mar 2026 v2

@@ 7 Recent Developments in Image Generation
Many top image generation systems have shifted from classic diffusion based formulations and moved toward Rectified Flow and Flow Matching, motivated by cleaner continuous time dynamics, improved training stability at scale, and image generation in fewer sampling steps. Moreover, the overall system has become more complicated and now consists of many pieces, including a large language model providing strong text understanding and better instruction following, more capable text encoders for precise conditioning, and a strong image backbone (often a Transformer) that converts those rich semantics into visual structure and detail. In the remainder of this section, we focus on the modeling shift itself: Rectified Flow and Flow Matching. Recently, they have emerged as compelling alternatives to diffusion objectives. We review the core formulation, and the key training and sampling choices that make these methods work well in practice. We also highlight the most impactful recent results that demonstrate their advantages for high quality and efficient image generation.
7.1 Background
@@
$x_{t+1}=x_{t}+f\!\left(x_{t},\theta_{t}\right)$ (101)
@@ 7.2 Rectified Flow
$\frac{dZ_{t}}{dt}=v_{\theta}\!\bigl(Z_{t},t\bigr)$ (105)
@@
Later, Esser et al. [113] scaled Rectified Flow for high-resolution image generation and were able to achieve superior performance compared to established diffusion models. While their training objective follows Rectified Flow, for the text to image architecture they use a multimodal diffusion backbone, DiT [102]. Fig. 70 shows the overall architecture of their model. With the Rectified Flow training setup, they are able to achieve competitive results and outperform the best open-source models and DALL-E 3 on GenEval [114]. Fig. 71 shows some of the sample images generated by their model with the corresponding prompts.
@@ 7.3 Flow Matching
with boundary conditions $\mu_{0}=0$, $\sigma_{0}=1$ at $t=0$ and $\mu_{1}=x_{1}$, $\sigma_{1}=\sigma_{\min}$ at $t=1$. They define an affine conditional map $\psi_{t}(\cdot)$ so that it pushes a standard Gaussian to a conditional Gaussian $p_{t}(x\mid x_{1})$
@@
where $\frac{d}{dt}\psi_{t}(z_{0};x_{1})$ is computed analytically from the definition of $\psi_{t}$. The authors show that, depending on the choice of the Gaussian path (i.e., $\mu_{t},\sigma_{t}$), the resulting probability path can correspond to either a diffusion path or an optimal transport path. The simplified training and inference algorithm for FM is shown in Box. 11.
They compare results using the same model architecture but trained with different objectives: one trained with optimal transport via Flow Matching, and another trained with a diffusion objective (DDPM [83]). As shown in Table XI, Flow Matching beats the diffusion based model in likelihood (BPD), quality of generated samples (FID), and evaluation time (NFE).
@@ 7.4 Conclusion
Box 11 A simplified overview of Flow Matching training (via regression to a target velocity along a chosen conditional path) and inference (via an Euler sampler).
@@ 8.1 Video Generation with GANs
Figure 73: Sample generated video frames from VGAN showing people walking on a golf course. Image source: [119]
@@
$\hat{x}^{(k)}_{I}=G_{I}\left(z_{I}^{(k)}\right),\qquad\hat{x}^{(k)}_{I}\in\mathbb{R}^{64\times 64\times 3}$ (116)
@@
The $D_{I}$ terms are the losses for the image discriminator, and the $D_{V}$ terms are the losses for the video discriminator. $R_{M}$ denotes the RNN parameters responsible for generating the motion codes $z_{M}$. $S_{1}$ and $S_{T}$ are random samplers for selecting 1 and $T$ consecutive samples, respectively. Using this approach, they produce video outputs that score a much higher preference by the user, as shown in Table XII.
@@ 8.3 Diffusion Based Video Generation
Using this approach, they are able to dramatically increase the inception score for unconditional video generation on the UCF-101 dataset, as shown in Table XIII. Fig. 77 shows text-conditioned video samples by cascading two models. First, samples are generated using a lower-resolution $16\times 64\times 64$ model. Then the samples are passed to a second model for simultaneous super-resolution and autoregressive extension to $64\times 128\times 128$.
@@ 8.4 Scaling up Video Diffusion Models
Blattmann et al. [131] introduced one of the first widely used open-source latent video diffusion models and called it Stable Video Diffusion (SVD). Most works prior to SVD only trained the video generation models on a small video dataset. Moreover, prior works mostly focused on temporal architectural changes rather than data and training strategy. The training pipeline consists of three stages. In stage I, they perform image pretraining with text–image pairs. Their experiments show that humans strongly prefer image pretrained models. In stage II, they pretrain the base video model. They preprocess a video clip and select $T$ frames and resize each frame to $H\times W=320\times 576$. They build conditioning by passing the text caption of the input video to a text encoder. The frame rate information is also embedded into a vector. The conditioning embeddings are then injected into the UNet block. The UNet block consists of both spatial (height, width) and temporal (frame) attention layers. They utilize the Elucidated Diffusion Model (EDM) [100] loss for training the UNet backbone.
Stage III has the same pipeline as stage II but they use 1 million high-quality text-to-video pairs that consist of lots of object motion, clean steady camera paths, and good captions. The resolution for this stage is $576\times 1024$ compared to $320\times 576$ at stage II. Box 12 shows their overall training pipeline. The authors also train an image-to-video variant of the model. For those, they start from the same base UNet and replace the text embedding with CLIP image embeddings and a few more tweaks. Fig. 78 shows some of the sample video frames generated by Stable Video Diffusion for different input text prompts.
@@ 9.4 DeepFake Detection and Benchmarks
Nataraj et al. [155] shifted towards an end-to-end framework to detect GAN-generated images, as shown in Fig. 90. They no longer rely on handcrafted features and instead let the CNN learn discriminative patterns from co-occurrence matrices computed separately on the R, G, and B channels. They evaluated the performance of the model on two different datasets: CycleGAN [162] and StarGAN [163]. CycleGAN contains image-to-image style transfer samples, and StarGAN consists of images of real celebrities, as well as GAN-generated celebrity images. In order to evaluate how well the model generalizes, they train on one dataset and use the other dataset for testing. Table XV shows the performance of the model on the CycleGAN and StarGAN datasets.
@@ 9.6 Watermarking Generated Images
As image generation models improved with the introduction of diffusion-based models, it became increasingly difficult to detect images generated by diffusion models. Wang et al. [166] ask what happens if they take the image through the inversion and reconstruction path. For diffusion-generated images, there should be a noise path that leads to a better reconstruction. In order to showcase this, they first take the image through the DDIM inversion process:
@@
$\mathrm{DIRE}(x_{0})=\lVert x_{0}-x_{0}^{\prime}\rVert$ (132)
where $x_{0}$ is the original image and $x_{0}^{\prime}$ is the reconstructed image. Fig. 94 shows the DIRE representation of real and diffusion-generated images. In experiments, real images show larger DIRE values compared to diffusion-generated ones. Lastly, they train a binary (ResNet-based) classifier on DIRE to detect diffusion-generated images. Using this approach, they are able to achieve better accuracy and average precision compared to other state-of-the-art detector models.
@@
where $\lambda_{i}$ is a weight on the image similarity loss. Note that only the signed decoder parameters of $D_{m}$ are updated but the encoder $E$ and extractor $W$ are kept frozen. Using a 48-bit signature, they are able to detect $99\%$ of newly generated images on the MS-COCO dataset while falsely flagging real images as fake only about one in $10^{9}$. Compared to Corvi et al. [168], they achieve an order of magnitude lower false positive rate for a similar true positive rate.
@@ 9.7 Conclusion
Image generation models have made significant leaps over the past decade. We have moved from models that produced low-quality, smudged outputs to models that generate images and coherent videos that are nearly indistinguishable from reality. Recent empirical scaling trends suggest continued improvements in the capabilities of these models. A central challenge is establishing effective safeguards against potential misuses. Do we have adequate safety measures in place? Are we prepared for a future in which the harmful uses of image generation systems, described above, are realized at scale? Given the risks associated with these powerful image and video generation models, we need both technical and societal solutions to mitigate these risks and to ensure preparedness for worst case scenarios.
@@ 10 Final Conclusion
Acknowledgement
We would like to thank Sean McGregor for his help with the publication.